Sol Summary – November 22th

Sol: 9
Summary Title: Trouble on Mars
Author’s name: Jas Purewal
Mission Status: Nominal
Sol Activity Summary:
A full day of EVA’s today. Morning walking search and rescue EVA. Afternoon EVA to test the robotic arm and collect biological samples. Crew Biologist is working on documentation of samples. Collected 43 unique individual samples.
Look Ahead Plan: XO, HSO and Biologist EVA at 10:30am to the Special Region, making a stop at Gateway to Lith for a Biological sample collecting mission. Afternoon will be Thanksgiving celebrations.
Anomalies in work: None
Weather: Rather nice. Sunny, clear. No rain.
Crew Physical Status: One crew member still feels a little dizzy, but improved from yesterday.
EVA: EVA 10 – XO and HSO walked to the Gateway to Candor and simulated a twisted ankle. EVA crew radioed into HABCOM for help. HABCOM sent two crew in two rovers (Opportunity and Curiosity) to rescue the injured crew member. All crew returned safely to the hab at the expected EVA end time. A very successful rescue mission was completed. The second EVA of the day, EVA 11 in the afternoon, was to Cowboy Corner. The Commander, Biologist and Engineer took Spirit and Perseverance. 7 samples collected including ice, root soil and plants. Pleiades was 50% successfully collecting rocks, it broke as expected from the handle and the pads from the long arms. These were the weak points that were deliberately pushed to breaking point. As it was breaking, it was visually confirmed that the operation was being affected by the weather (e.g. accumulation of debris).
Reports to be filed:
Commander’s report
EVA report
EVA request
Operations report
Sol Summary report
Journalist report

Support Requested: none

Copyright © The Mars Society. All rights reserved. | Main Site